OpenStudy (anonymous):

Complete the equation that is related to the given equation. t – 1.8 = –2.3 t = –2.3 +

OpenStudy (anonymous):

wat do u think

OpenStudy (anonymous):

-4.1???

OpenStudy (anonymous):

how did you get dat

OpenStudy (anonymous):

internet coculator

OpenStudy (anonymous):

my spellimg is terrible iknow

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i got -0.5

OpenStudy (firekat97):

I'm pretty sure its 1.8 because t - 1.8 = -2.3 t -1.8 + 1.8= -2.3 + 1.8 simplify left hand side, t = -2.3 + 1.8 and we are told that t = -2.3 + (x) therefore (x) = 1.8 does that make sense?